### v2

# ReLoop: "Seeing Twice and Thinking Backwards" via Closed-loop Training to Mitigate Hallucinations in Multimodal understanding

Abstract
While Multimodal Large Language Models (MLLMs) have achieved remarkable progress in open-ended visual question answering, they remain vulnerable to hallucinations. These are outputs that contradict or misrepresent input semantics, posing a critical challenge to the reliability and factual consistency. Existing methods often rely on external verification or post-hoc correction, lacking an internal mechanism to validate outputs directly during training. To bridge this gap, we propose ReLoop, a unified closed-loop training framework that encourages multimodal consistency for cross-modal understanding in MLLMs. ReLoop adopts a ring-shaped structure that integrates three complementary consistency feedback mechanisms, obliging MLLMs to "seeing twice and thinking backwards". Specifically, ReLoop employs the frozen Consistency Feedback Plugin (CFP), comprising semantic reconstruction and visual description modules, along with an attention supervision module for attention alignment. These components collectively enforce semantic reversibility, visual consistency, and interpretable attention, enabling the model to correct its outputs during training. Extensive evaluations and analyses demonstrate the effectiveness of ReLoop in reducing hallucination rates across multiple benchmarks, establishing a robust method for hallucination mitigation in MLLMs. The code is available at: https://github.com/ZiyanHuang11/Reloop-hallucinations.
ReLoop: "Seeing Twice and Thinking Backwards" via Closed-loop Training to Mitigate Hallucinations in Multimodal understanding
Jianjiang Yang1thanks: Equal contribution, Yanshu Li211footnotemark: 1, Ziyan Huang311footnotemark: 1 1University of Bristol, 2Brown University, 3South China University of Technology edisonyang109@gmail.com, yanshu_li1@brown.edu, bonnie.ziyan.huang@gmail.com
1 Introduction
Figure 1: Illustration of four major hallucination types in open-ended VQA. Despite being visually grounded, MLLMs produce fluent but hallucinated responses across object, attribute, relation, and event dimensions.
In recent years, MLLMs (Liu et al., 2023b; OpenAI, 2023; Li et al., 2023a) have demonstrated significant progress in bridging vision and language, addressing tasks such as visual question answering (VQA), image captioning, and instruction adherence. However, a fundamental difficulty that persists is hallucination, where the generation of outputs that are inconsistent with or unsupported by visual inputs (Kalavasis et al., 2024). Hallucinations are especially common in open-ended VQA circumstances, in which unclear or underspecified questions can result in factual mistakes. These hallucinations span diverse categories, including Object, Attribute, Relation, and Event. Figure 1 illustrates that a singular image of "a dog grasping an orange frisbee" can elicit various forms of hallucination: a fictitious "cat" (object), an incorrectly identified "green" frisbee (attribute), an erroneous spatial relation "under the paw" (relation), or a temporal misrepresentation "sleeping" (event). These errors are semantically plausible yet visually unfounded, posing major challenges for trustworthiness and safety of MLLMs across critical applications, including medical decision-making (Kim et al., 2025), robotic perception (Park et al., 2023), and autonomous navigation (Alsulaimawi, 2025).
Existing works  (Sun et al., 2023; Ayala and Béchard, 2024; Sun et al., 2024) often regard hallucination as an output-level anomaly that is corrected post hoc, overlooking its underlying cause. In practice, hallucinations frequently arise from misalignment between the input, visual content, and the model’s latent reasoning. Without an internal supervision mechanism, models may produce fluent yet ungrounded answers. We argue that hallucination stems from the model’s inability to validate its own output across modalities and recommend injecting this ability directly into training.
We subsequently derive inspiration from human cognitive processes. When answering visual questions, individuals rarely rely on a single forward guess. Instead, after answering, they may reassess the question’s intent, examine the visual scene, and refine conclusions—especially in the face of ambiguity or uncertainty. However, most models operate in a unidirectional manner, mapping $(Q,I\rightarrow A)$. As a result, once the model makes a prediction, there is no structured way to assess whether it actually understood the question, if the answer aligns with the visual evidence, or whether the model attended to the right regions in the image.
To address this issue, we propose ReLoop, a cognitively inspired unified training framework that encourages multimodal consistency for cross-modal understanding in MLLMs. ReLoop implements a feedback-driven closed-loop supervision process, allowing the model to reassess its predictions and validate their consistency with the original input through multi-level supervision during training. Specifically, after MLLMs produce an answer from the image-question pair, Reloop enables the model to recapture input semantics and assess internal consistency via: a Consistency Feedback Plugin (CFP), comprising two frozen modules: (1) CFP-Lang reconstructs the question $\hat{Q}^{*}$ from $(A,I)$ to supervise semantic alignment, and (2) CFP-Vis generates a description ${I}^{*}$ to assess factual grounding. In parallel, an attention supervision module extracts the model’s token-to-image attention map $\mathcal{H}$ and compares it with an entropy-based pseudo-ground truth. All signals are integrated as differentiable losses in the overall optimization objective. This design encourages the model to "see twice and think backward"—first look to answer ($Q,I\rightarrow A$), look twice to reassess ($A,I\rightarrow\hat{Q}^{*},{I}^{*},\mathcal{H}$), and finally to correct ($\hat{Q}^{*},{I}^{*},\mathcal{H}$ $\approx Q,I,\mathcal{H}_{\text{pseudo}}$).
ReLoop bridges the gap between perception and output. It turns the black-box understanding process into an interpretable, feedback-aware loop that continuously refines the model’s internal representations. Our key contributions can be summarized clearly as follows:
  • We propose ReLoop, a cognitively inspired closed-loop training framework that ensures consistency among image, question, and answer modalities, effectively mitigating hallucinations in MLLMs.
  • We introduce three complementary consistency signals: semantic reconstruction, visual description, and attention alignment, to emulate the humanlike "reversible thinking" process and improve cross-modal consistency during training.
  • We provide a novel use of pretrained vision-language models by repositioning them as frozen Consistency Feedback Plugins (CFPs) in the training loop. Rather than functioning as typical forward-only encoders, they now perform in a reflective, backward supervisory role, producing feedback signals to guide the main model’s alignment with multimodal semantics.
Figure 2: Seeing Twice and Thinking Backwards: ReLooping Hallucination Suppression in Multimodal Language Models. This diagram aligns human cognitive phases (left) with model modules (right) in a closed-loop process. The main model $M$ produces an answer which is then introspected via CFP‑Lang (language reconstruction), CFP‑Vis (visual description), and internal cross‑attention maps. Semantic aggregation, CLIP similarity, and entropy‑based soft masks produce feedback losses that are summed and back‑propagated to update $M$ and the semantic aggregator $S$.
2 Related Work
Hallucination Mitigation in MLLMs.
Multimodal LLMs frequently produce hallucinations: responses conflicting with visual inputs, such as inventing entities or misaligning semantics (Li et al., 2023b). Recent mitigation efforts combine post-hoc correction and architectural refinement. Retrieval-augmented methods like (Mala et al., 2025) grounds outputs in external knowledge via hybrid retrievers, while (Ayala and Béchard, 2024) reduces hallucinations in structured outputs. Architectural solutions such as OPERA (Huang et al., 2024) penalize over-trust during decoding, and preference-aligned training like TPO (Gu and Wang, 2025) enhances vision grounding. Post-generation verification (Woodpecker (Yin et al., 2023)) and decoding-time personalization (PAD (Chen et al., 2024)) complement training-time alignment; concurrent work improves multimodal ICL efficiency and task control (Li et al., 2025a, b) and instruction-tuned dialog grounding (Luo et al., 2024b). Beyond images, video hallucination is diagnosed via fine-grained spatio–temporal grounding (Luo et al., 2025).
Semantic Reversibility and Bidirectional Supervision.
Human cognition leverages bidirectional reasoning to validate hypotheses: a principle termed "cognitive reversibility" (Johnson-Laird, 1983). Recent works explore this idea through decoding-time strategies: Self-RAG (Asai et al., 2023) integrates retrieval-augmented generation with self-reflection, enabling models to critique and refine their outputs iteratively, while DeepSeekMath employs Group Relative Policy Optimization (GRPO) (Shao et al., 2024), enhancing mathematical reasoning by optimizing policy decisions based on group sampling strategies. Similarly, back-translation methods (Sennrich et al., 2016) enforce answer-question consistency through round-trip translation.
Cross-modal Consistency.
Ensuring cross-modal consistency is vital for mitigating hallucinations in MLLMs. Recent methods enhance visual-text alignment to reduce semantic drift. VCD (Leng et al., 2024) contrasts outputs from original and perturbed images to promote grounding and reduce unimodal bias. HACL (Jiang et al., 2024) treats hallucinated captions as hard negatives to improve alignment. EAGLE (Villa et al., 2025) further refines visual encoders post-pretraining, yielding better grounding and fewer hallucinations. Broader benchmark efforts catalog the MLLM evaluation landscape (Li et al., 2024), including conversational aspect-based sentiment settings (Luo et al., 2024a).
3 Preliminaries
3.1 Task Formulation: Open-ended Visual Question Answering
We consider the task of open-ended VQA, where the model receives an image $I$ and a natural language question $Q$, and produces a free-form answer $A$. Unlike multiple-choice settings, this task requires the model to produce linguistically coherent and visually grounded responses without predefined options.
In this case, hallucination refers to answers that contradict the image $I$, misinterpret the question $Q$, or introduce unsupported content.
3.2 Consistency Signals
To encourage faithful understanding, we supervise the model using three types of cross-modal consistency signals:
Linguistic Consistency. We verify whether the model’s answer $A$ implies the same question intent as the original $Q$, by attempting to reconstruct $Q$ from $(A,I)$. This tests whether the model understood the question meaningfully.
Visual Consistency. We evaluate whether the answer $A$ is factually grounded in image $I$, by generating a descriptive caption ${I}^{*}$ based on $(A,I)$ and checking its alignment with the image. This ensures that the response reflects the actual visual content.
Attention Consistency. We examine whether the model attends to the correct regions of the image while producing $A$. This is assessed by comparing its internal attention map $\mathcal{H}$ with a soft pseudo-ground truth $\mathcal{H}_{\text{pseudo}}$ derived from entropy-based cues.
Together, these consistency signals serve as indirect evidence of whether the model truly grasps both the visual input and the question semantics.
4 ReLoop Framework: Reflect, Recapture, and Optimize through a Closed-Loop Process
We introduce ReLoop, a unified training framework aimed at reducing hallucinations in MLLMs for open-ended VQA answering. As illustrated in Figure 2, the framework incorporates three complementary consistency feedback mechanisms: semantic reconestruction, visual description, and attention alignment to supervise the model toward producing answers faithful to both the question and the image.
These feedback signals are instantiated through a frozen Consistency Feedback Plugin (CFP): semantic reconstruction (CFP-Lang) and visual description (CFP-Vis), and attention supervision from the model itself. The CFP module is broadly compatible with a range of encoder-decoder or decoder-only MLLMs. During inference (First Look → Answer), the model receives a question-image pair and produces an initial answer. The training process then begins with Reflect → Second Look → Correct: the model examines its output through structured consistency feedback. Specifically, it "introspectively" asks:
  • "Did I understand the QQ?" (→ semantic reconstruction)
  • "Did I say it right?" (→ visual description)
  • "Did I focus on the right region?" (→ attention alignment)
ReLoop decomposes hallucination mitigation into two interacting components:
  • "Re" emphasizes recapturing details, encouraging the model to reassess the semantic and visual cues from both question and image through CFP modules and token-level attention heatmaps.
  • "Loop" denotes a feedback-driven training loop. After each forward prediction, feedback from the three consistency pathways is aggregated into the loss function ($L_{\text{align}}$, $L_{\text{vis}}$, $L_{\text{attn}}$), driving iterative updates that refine the model’s multimodal grounding and answer reliability.
4.1 A Closed-loop Training
The entire training process follows a closed-loop pattern, emulating "seeing twice and thinking backward". Each training step proceeds as follows:
  1. 1.
    First Look: The main model $M$ takes the image $I$ and question $Q$ as input to produce an initial answer $A$.
  2. 2.
    Reflect: The model introspects on $A$ by reconstructing a proxy question $\hat{Q}$, generating a visual description ${I}^{*}$, and extracting token-level attention $\mathcal{H}$.
  3. 3.
    Second Look: The reconstructions are compared against the original inputs to compute consistency losses, capturing discrepancies in semantics, visual grounding, and attention focus.
  4. 4.
    Correct: All feedback signals are aggregated into $L_{\text{total}}$ to update $M$ and the semantic aggregator $S$ via backpropagation.
This multi-stage loop is repeated across training epochs, leading to the model $M$ that gradually reduces hallucinations.
4.2 Re: Recapturing Details for Consistency Supervision
This stage corresponds to the training-time processes of "Reflect" and "Second Look", where the model reassesses its answers to recapture overlooked semantic and visual details. Three feedback pathways modules examine whether the model understood the question, correctly grounded its answer in the image, and attended to salient regions.
Figure 3: Token-Level Attention Supervision. Visualization of predicted attention $\mathcal{H}$ and entropy-based pseudo ground truth $\mathcal{H}_{\text{pseudo}}$ for two key answer tokens: dog (top row) and playing (bottom row).
4.2.1 CFP-lang: Language Reconstruction and Adaptive Consistency Weighting
To evaluate whether the model correctly interprets the input question, we introduce a frozen language reconstruction module, CFP-lang. Given the answer-image pair $(A,I)$, CFP-lang produces a set of candidate reverse questions $\{\hat{Q}_{1},\hat{Q}_{2},\ldots,\hat{Q}_{k}\}$ that approximate possible intents underlying the predicted answer. A lightweight semantic aggregator $S$, composed of a BERT encoder and a single-layer MLP, scores each candidate against the original question $Q$ using BERTScore. The highest-ranked proxy $\hat{Q}^{*}$ is selected to reflect the model’s inferred intent.
However, directly enforcing alignment on all reconstructed questions may introduce noise, particularly when the produced answer is short or underspecified. To mitigate this, we introduce an Adaptive Consistency Weighting (ACW) mechanism, which adjusts the attention supervision (mentioned in section 4.2.3) strength based on the similarity between $Q$ and $\hat{Q}^{*}$:
$\gamma=\begin{cases}1.0&\text{if BERTScore}(Q,\hat{Q}^{*})\geq 0.8\\ 0.1&\text{if }0.6\leq\text{BERTScore}(Q,\hat{Q}^{*})<0.8\\ 0.01&\text{if BERTScore}(Q,\hat{Q}^{*})<0.6\end{cases}$ (1)
Rather than discarding low-confidence pairs, this soft weighting ensures that stronger semantic matches contribute more prominently to the learning objective. The language consistency loss is computed as:
$L_{\text{align}}=1-\text{BERTScore}(Q,\hat{Q}^{*})$ (2)
4.2.2 CFP-visual: Visual Description and Similarity Supervision
To validate whether the produced answer $A$ is visually grounded in the image $I$, we employ a frozen visual description module, CFP-visual. Given $(A,I)$, it generates a caption ${I}^{*}$ describing the image content implied by the answer. We then compute the cosine similarity between the CLIP-encoded vectors of $I$ and ${I}^{*}$, and derive the visual consistency loss as:
$L_{\text{vis}}=1-\cos(\text{CLIP}_{\text{img}}(I),\text{CLIP}_{\text{text}}({I}^{*}))$ (3)
4.2.3 Attention Supervision via Heatmap Consistency
To enhance interpretability and mitigate hallucinations arising from inattentive or unstable decoding, we explicitly supervise the model’s token-level cross-attention patterns. From the decoder of the main model $M$, we extract attention maps $\mathcal{H}$, which indicate the spatial focus during answer generation. We construct a soft pseudo-ground-truth heatmap $\mathcal{H}_{\text{pseudo}}$ using entropy-based masking (Detailed explanation can be found in Appendix C). This method preserves uncertainty information and avoids brittle hard labels. As illustrated in Figure 3, well-grounded tokens (e.g., dog) yield concentrated heatmaps aligned with visual evidence, while hallucinated tokens (e.g., playing) produce offset patterns.
[…]
Object Visual CLIP(II, ${I}^{*}$) 28.02 $\pm$ 3.10 29.46 $\pm$ 3.27 $\uparrow$1.44 24.5% 10.3% $\downarrow$14.2%
Language BERT($Q$, $\hat{Q}$) 0.862 $\pm$ 0.022 0.873 $\pm$ 0.024 $\uparrow$0.011
Attention Entropy($\mathcal{H}$) 1.31 $\pm$ 0.40 1.28 $\pm$ 0.45 $\downarrow$0.03
Attribute Visual CLIP(II, ${I}^{*}$) 26.59 $\pm$ 3.31 26.81 $\pm$ 3.41 $\uparrow$0.22 7.3% 4.0% $\downarrow$3.3%
Language BERT($Q$, $\hat{Q}$) 0.868 $\pm$ 0.025 0.894 $\pm$ 0.028 $\uparrow$0.026
Attention Entropy($\mathcal{H}$) 1.36 $\pm$ 0.46 1.32 $\pm$ 0.52 $\downarrow$0.04
Relation Visual CLIP(II, ${I}^{*}$) 27.22 $\pm$ 3.26 28.01 $\pm$ 3.38 $\uparrow$0.79 13.2% 7.6% $\downarrow$5.6%
Language BERT($Q$, $\hat{Q}$) 0.855 $\pm$ 0.020 0.875 $\pm$ 0.023 $\uparrow$0.020
Attention Entropy($\mathcal{H}$) 1.39 $\pm$ 0.43 1.34 $\pm$ 0.50 $\downarrow$0.05
Event Visual CLIP(II, ${I}^{*}$) 26.63 $\pm$ 3.08 26.94 $\pm$ 3.37 $\uparrow$0.31 10.4% 5.2% $\downarrow$5.2%
Language BERT($Q$, $\hat{Q}$) 0.861 $\pm$ 0.024 0.877 $\pm$ 0.029 $\uparrow$0.016
Attention Entropy($\mathcal{H}$) 1.33 $\pm$ 0.42 1.51 $\pm$ 0.55 $\uparrow$0.18
Table 1: Effect of ReLoop on consistency and hallucination reduction across different hallucination types. We compare MiniGPT-4 (baseline) and ReLoop in terms of signal outputs from three frozen feedback modules: visual grounding (CLIP similarity), semantic alignment (BERTScore), and attention focus (entropy). $\Delta$ denotes the absolute change in signal quality after applying ReLoop.
We enforce alignment between $\mathcal{H}$ and $\mathcal{H}_{\text{pseudo}}$ by minimizing the KL divergence:
$L_{\text{attn}}=\text{KL}(\mathcal{H}\parallel\mathcal{H}_{\text{pseudo}})$ (4)
4.3 Loop: Feedback Aggregation, Alignment, and Optimization
After consistency signals are computed from language, vision, and attention supervision, ReLoop aggregates them into a unified training objective. This stage corresponds to the "Correction" step in the loop, where the model updates its parameters based on multi-perspective feedback. The total loss combines standard supervision with the three consistency terms:
$L_{\text{total}}=L_{\text{sft}}+\alpha\cdot L_{\text{align}}+\beta\cdot L_{\text{vis}}+\gamma\cdot L_{\text{attn}}+\lambda\cdot\Omega(\theta)$ (5)
where $L_{\text{sft}}$ is the token-level cross-entropy loss, and $\Omega(\theta)$ is an L2 regularization term. The consistency weights are empirically set as $\alpha=1.0$, $\beta=0.7$, $\lambda=10^{-5}$ and $\gamma$ is defined in Equation 1.
Only the parameters of the main model $M$ and the semantic aggregator $S$ are updated during training. All feedback modules, including CFP-Lang, CFP-Vis, attention supervision, and CLIP, remain frozen.
5 Experimental Setup
[…]
Figure 4: Type‑wise hallucination rates (%) for baseline (MiniGPT-4) and ReLoop models.
Evaluation Benchmarks and Metrics. We evaluate ReLoop on a broad range of hallucination and multimodal understanding benchmarks, including POPE (Li et al., 2023b), CHAIR (Rohrbach et al., 2018), AMBER (Wang et al., 2023), MMHal-B (Sun et al., 2023), HallusionBench (Guan et al., 2024), Faith/FaithS (Jing et al., 2024), and MME (Fu et al., 2023). Full definitions can be found in Appendix A.3 A.4.
Figure 5: KDE distributions of CLIP similarity, BERTScore, and attention entropy for hallucinated and non-hallucinated samples. ReLoop’s frozen modules exhibit sharp signal shifts that serve as reliable supervision sources.
Baselines. We use MiniGPT-4 as the baseline model in Experiment 6.1 and compare against LLaVA-1.5 variants trained with LLaVA-RLHF Sun et al. (2023), HA-DPO Zhao et al. (2023), POVID Zhou et al. (2024), and Visual Contrastive Decoding (VCD) Leng et al. (2023). For robustness analyses, we adopt LLaVA-1.5 + ReLoop as the canonical setting and report stress tests under noisy external supervision and nonsensical answers (Table 5). Unless otherwise specified, all baselines share the same backbone, data, and training protocol for a fair comparison. Implementation details are provided in Appendix A.5.
Model Hallucination Suppression Cross-modal Faithfulness
[…]
MiniGPT-4 82.3 49.0 22.7 63.2 86.7 68.5
 + ReLoop 83.9 38.8 20.5 69.9 88.6 71.3
InstructBLIP 83.8 47.8 20.6 68.4 87.3 69.8
 + ReLoop 85.3 36.9 17.5 67.0 88.5 73.2
LLaVA-1.5 85.7 53.5 24.2 65.8 89.5 75.8
 + ReLoop 86.3 40.2 16.2 70.3 89.2 75.3
LLaVA-1.6 86.8 52.0 21.8 67.3 89.4 76.6
 + ReLoop 87.9 38.5 16.1 71.1 89.2 76.2
Qwen-VL-2.5 89.3 47.3 20.8 69.1 89.7 76.5
 + ReLoop 90.7 37.6 16.6 72.5 90.4 77.8
mPLUG-owl 89.1 62.5 31.0 58.9 88.3 72.7
 + ReLoop 90.9 42.5 21.8 66.5 87.9 71.0
ShareGPT4V 88.2 50.2 21.8 68.0 88.2 73.6
 + ReLoop 89.7 44.9 21.5 69.2 89.3 74.8
Table 2: Performance comparison of various LVLMs with and without ReLoop. Hallucination is measured by POPE, CHAIRs, and CHAIRi; cross-modal faithfulness is evaluated using F1, Faith, and FaithS. ↓ indicates lower is better; ↑ indicates higher is better.
6 Results and Analysis
6.1 Identify Internal Causes of Hallucinations: Module Signals vs. Hallucination States
We first aim to pinpoint internal representation deficiencies that drive hallucination behaviors across different hallucination types. We analyze consistency signal deviations produced by ReLoop’s frozen supervision modules, with hallucinated versus non-hallucinated samples. Responding: "Did I understand the question?” (language, via BERTScore); "Did I say it right?” (visual, via CLIP similarity); "Did I focus on the right region?” (attention, via entropy).
Multimodal hallucinations stem from structured, modality-specific representation gaps. As shown in Table 1, hallucinated responses are consistently associated with lower CLIP similarity (–2.25), reduced BERTScore (–0.034), and higher attention entropy (+0.31). Figure 5 reveals distinct signal patterns associated with different hallucination types. Object hallucinations correspond to a clear leftward shift in CLIP similarity, indicating weaker visual grounding. Attribute hallucinations are marked by lower BERTScore, reflecting reduced semantic alignment. Event hallucinations show higher attention entropy, suggesting that the model distributes focus more broadly, which may help in capturing complex scenes but also increases the risk of focusing on irrelevant regions.
Signal dynamics vary by hallucination type. (1) Object hallucinations are primarily rooted in the visual module. They often manifest as hallucinated entities not present in the image. ReLoop yields a significant gain in CLIP similarity (↑1.44) and a decrease in attention entropy (↓0.03), suggesting enhanced image-text alignment and focused visual grounding. (2) Attribute hallucinations show the largest improvement in BERTScore (↑0.026) and only a slight change in CLIP similarity (↑0.22), indicating that semantic reconstruction plays a more important role than visual grounding. This aligns with their nature: attributes often relate to textual misinterpretation (e.g., color or size), even when visual cues are present. (3) Relation hallucinations involve complex spatial or relational semantics and display moderate improvements across all three signals (CLIP↑0.79, BERT↑0.020, Entropy↓0.05), suggesting that ReLoop’s multi-signal supervision addresses cross-modal misalignment collaboratively. (4) Event hallucinations are primarily tied to attention misallocation. ReLoop improves CLIP (↑0.31) and BERT (↑0.016) slightly, but entropy increases (↑0.18), reflecting broader attention scopes. This likely helps avoid fixation on irrelevant regions, especially in dynamic or temporally inferred scenes. Figure 4 shows that ReLoop successfully mitigates hallucinations compared to MiniGPT-4 across four hallucination types.
Ablation Version Hallucination Suppression Cross-modal Faithfulness
[…]
MiniGPT-4 83.0 49.0 22.7 60.2 84.3 64.2
w/o Consistency Supervision 84.2 47.4 21.6 60.7 86.7 68.5
w/o Gating & Aggregator 85.4 39.8 19.7 60.4 88.1 71.6
w/o Attention Supervision 83.6 40.2 20.1 61.9 86.3 67.5
Full ReLoop 84.9 38.3 18.9 63.1 88.6 72.8
Table 3: Performance comparison of ReLoop under different ablation configurations on MiniGPT-4. Removing consistency supervision results in the worst faithfulness and hallucination rate, while full ReLoop delivers the best overall performance. Although gating removal slightly improves POPE, it hurts precision (F1) and consistency.
Method Hallucination Suppression Cross-modal Faithfulness
[…]
LLaVA-1.5 83.5 53.9 23.5 63.2 86.9 70.5
+ LLaVA-RLHF 88.2 44.5 20.1 67.0 89.0 74.4
+ HA-DPO 86.7 52.3 21.6 65.4 88.4 73.5
+ POVID 84.3 53.2 24.2 64.7 87.3 71.8
+ VCD 86.8 43.1 20.2 66.9 88.8 73.6
+ ReLoop 87.9 42.0 19.5 67.4 89.5 75.1
Table 4: Performance comparison of ReLoop with alignment-enhancing baselines for LLaVA-1.5 on hallucination suppression and cross-modal faithfulness. Best scores are in bold and the second best are underlined.
6.2 Effects of Structured Feedback in ReLoop
Motivated by earlier findings, we evaluate how effectively ReLoop’s structured feedback enhances semantic grounding across five representative LVLMs (Table 2). The observed improvements span models with diverse architectures and training paradigms, showing that ReLoop is broadly compatible and easily integrable into various LVLMs.
[…]
Cross-modal Faithfulness. ReLoop also enhances cross-modal faithfulness. F1 increases on MiniGPT-4 (+10.6%), LLaVA-1.5 (+6.8%), LLaVA-1.6 ($\sim$5.6%), and Qwen-VL-2.5($\sim$4.9%); InstructBLIP maintains comparable F1 while gaining on faith metrics. FaithS improves for MiniGPT-4 (+2.8), InstructBLIP (+3.4), Qwen-VL-2.5 (+1.3), and ShareGPT4V (+1.2), and remains near-parity on LLaVA-1.6. These gains suggest that the model not only grounds responses more accurately in the image but also maintains semantic alignment with the question intent.
6.3 Robustness Under Noisy Supervision
We next ask whether closed-loop training remains stable when external supervision is imperfect or when the initial answer signal is degenerate. We stress-test ReLoop under (i) noisy teacher feedback on the visual channel and (ii) nonsensical answers that could mislead the loop.
[…]
LLaVA-1.5 + ReLoop (clean) 87.9 42.0 19.5 52/41/9
+ Teacher noise (15%) 86.8 42.6 20.2 0.11-0.11
+ Answer noise (15%) 87.2 43.1 20.4 35/43/22 0.06-0.06
Table 5: Robustness under noisy supervision. Teacher-side visual description corruption (15%) and answer-side nonsense injection (15%) have limited impact on core hallucination metrics. CLIP similarity drops under teacher noise, whereas ACW re-allocates per-sample weights under answer noise (smaller high-confidence mass).
[…]
Natural noise robustness via fourfold filtering. ReLoop’s robustness emerges from a fourfold filter: three orthogonal supervision signals (language via BERTScore, visual via CLIP, attention via entropy-aware $\mathcal{H}_{\text{pseudo}}$) plus ACW’s discrete gating $\gamma\!\in\!\{1,0.1,0.01\}$ (Sec. 4.2). Noisy teacher feedback is first cross-validated across modalities, then attenuated by ACW, and finally diluted in the multi-loss objective (Sec. 4.3), so that biased cues have limited influence on the update direction. Empirically, the stability of POPE/CHAIR under teacher-side corruption and the expected shift in the $\gamma$ distribution under answer nonsense together indicate that the closed loop self-regularizes and converges stably despite imperfect teachers.
6.4 Ablation Study
To assess the contribution of each component in ReLoop, we perform a coarse-grained ablation study over four configurations (Table 3). Removing consistency supervision leads to the highest hallucination rates (CHAIRs: 47.4) and lowest semantic faithfulness (FaithS: 68.5), highlighting its central role. Attention supervision also proves important, as its removal moderately reduces FaithS. While removing gating slightly improves POPE, it harms F1 and hallucination suppression. Full ReLoop achieves the best overall results, reducing CHAIRs by 10.7 and increasing FaithS by 8.6 over the baseline. These findings underscore the complementary roles of all modules and the importance of structured feedback for robust alignment.
6.5 Unified Comparison with Alignment Strategies
We compare ReLoop with representative alignment methods, LLaVA-RLHF, HA-DPO, and POVID on both fine-grained hallucination metrics and broader benchmark evaluations. As shown in Table 4, ReLoop consistently
[…]
LLaVA-1.5 73.9 1513 65.4 48.6
+ LLaVA-RLHF 73.8 1231 64.3 43.2
+ HA-DPO 77.2 1374 65.6 49.9
+ POVID 75.8 1421 65.9 51.4
+ ReLoop 80.3 1505 68.9 52.3
Table 6: Benchmark-level comparison of ReLoop with alignment strategies across four evaluation baselines.
outperforms alternatives on POPE, CHAIR, F1, and faithfulness metrics, indicating stronger hallucination suppression and cross-modal faithfulness. On benchmark-level evaluations (Table 6), ReLoop leads on AMBER, MMHal-B, and HallusionBench, while remaining competitive on MME. The slight MME drop may reflect a common trade-off between alignment supervision and low-level perception, also observed in other alignment-based methods like LLaVA-RLHF. These findings underscore ReLoop’s effectiveness across both targeted and comprehensive settings.
6.6 Additional Analyses
To further substantiate the effectiveness and practicality of ReLoop, we provide supplementary analyses in the appendices. Appendix A.6 reports a Training Cost Breakdown: cross-method cost, CFP overhead attribution, and convergence/epoch-level timing. Appendix A.7 ablates contrastive augmentation to isolate closed-loop gains. Appendix B provides a Case Study over four hallucination types and a nonsensical-answer failure mode, illustrating early rejection, ACW $\gamma$-downweighting, and entropy-aware masking.
7 Conclusion
We present ReLoop, a closed-loop training framework that mitigates hallucinations in MLLMs by enforcing semantic and visual consistency through bidirectional feedback. By incorporating language reconstruction, visual description, and attention alignment, ReLoop allows models to verify and refine predictions during training. Experiments show consistent gains in hallucination suppression and interpretability, establishing ReLoop as a general foundation for building more reliable MLLMs.
Potential Limitations
Performance Variability Across Hallucination Types. While ReLoop substantially improves hallucination suppression in object and attribute categories, its effectiveness on relation and event hallucinations remains relatively modest. These hallucination types often involve higher-order reasoning and temporal or spatial understanding, which are less easily corrected through current consistency signals. Future extensions may incorporate specialized supervision tailored to relational semantics or causal cues to address this gap.
Supervision Dependency and Domain Adaptability. ReLoop relies on access to paired image–question–answer data to compute consistency signals. This requirement poses challenges in domains with limited high-quality supervision, such as medical or scientific imaging. Moreover, the training framework assumes reasonably clean and grounded reference answers, which may not hold in low-resource or noisy environments. Reducing ReLoop’s dependence on strongly supervised inputs and exploring semi-supervised or synthetic feedback generation remain promising directions for broader applicability.
Scope of evaluation. Our study focuses on standard VQA-style image benchmarks and general-domain LVLMs. We have not evaluated text-heavy or long-tail domains (e.g., dense OCR, charts) or temporal reasoning tasks, where attention allocation and signal reliability may differ. Extending evaluation to these regimes is left for future work.
Ethics Statement
All datasets utilized in this work are either publicly released or ethically sourced, ensuring full compliance with associated data usage policies. For evaluation purposes, we additionally include AI-generated content produced under controlled prompting conditions. These samples are clearly labeled and subjected to careful human verification to ensure factual accuracy and annotation quality. We acknowledge the broader implications of hallucination mitigation in AI systems and advocate for responsible model development that prioritizes reliability, fairness, and interpretability.
Appendix A Additional Experimental Details
A.1 Implementation Details
Backbone and Setup.
We apply ReLoop to five representative LVLMs with diverse architectures: MiniGPT-4, InstructBLIP, LLaVA-1.5, mPLUG-owl, and ShareGPT4V. To assess generalizability on stronger backbones, we further evaluate on LLaVA-1.6 and Qwen-VL-2.5 (Table 2). Importantly, we do not alter the internal structures of these models. ReLoop is introduced as a lightweight, external consistency-supervision framework during training. All backbones are initialized with their public checkpoints and keep their visual encoders (e.g., ViT, CLIP) frozen.
ReLoop Components.
ReLoop introduces three frozen feedback modules: (1) CFP-Lang (MiniGPT-4-based reverse question reconstructor); (2) CFP-Vis (BLIP-2-based visual describer); (3) Attention Supervision that aligns decoder attention maps with entropy-based soft pseudo-labels. A frozen BERT encoder plus an MLP scorer serves as a lightweight semantic aggregator. All feedback modules remain frozen; only the backbone and the aggregator are updated.
Training Details.
[…]
Loss Function.
The overall objective is
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{sft}}+\alpha\,\mathcal{L}_{\text{align}}+\beta\,\mathcal{L}_{\text{vis}}+\gamma\,\mathcal{L}_{\text{attn}}+\lambda\,\Omega(\theta)$ (6)
We set the hyper-parameters as $\alpha=1.0$, $\beta=0.7$, and $\lambda=10^{-5}$. The weight $\gamma$ is dynamically adjusted by the Adaptive Consistency Weighting (ACW) mechanism, which modulates $\gamma$ based on the BERTScore between the original and reconstructed questions (see Section 4.2.1).
A.2 Training Dataset Construction
We curated approximately 30K high-quality QA-image triplets from the LLaVA-Instruct-150K corpus (Liu et al., 2023a), each containing an image, an open-ended question, and a human-annotated answer. To simulate hallucination supervision, we generated semantically contradictory answers by modifying key elements (e.g., objects, attributes, or relations) in the references. These hallucinated samples were automatically constructed and manually verified for quality and type diversity. In Experiment 6.1, we selected 500 representative QA-image pairs from the filtered validation set based on POPE and MMHalBench, equally split between hallucinated and non-hallucinated cases. In Experiment 6.2, we evaluated five LVLMs on this curated set to assess the impact of ReLoop. Models with open alignment architectures (e.g., MiniGPT-4, InstructBLIP) showed the greatest improvement, while high-performing black-box models (e.g., ShareGPT4V) saw minimal gains, suggesting ReLoop’s effectiveness hinges on alignment signal compatibility.
A.3 Evaluation Metrics
To comprehensively evaluate the effectiveness of ReLoop in mitigating hallucinations and enhancing visual grounding, we adopt a structured set of metrics covering both hallucination suppression and cross-modal consistency. In particular, shown in Table 4, we group the metrics into two key categories: Hallucination Suppression, which quantifies the presence of non-existent or spurious content, and Cross-modal Faithfulness, which assesses the semantic and perceptual alignment between generated text and visual input.
A.3.1 Metrics on Hallucination Suppression
For hallucination evaluation, we incorporate CHAIR Rohrbach et al. (2018) to measure hallucination frequencies at instance levels and include POPE Li et al. (2023b), a probing-based diagnostic benchmark to evaluate object hallucinations through direct VQA-style interactions. Together, these metrics allow us to holistically assess ReLoop’s ability to suppress hallucinated content while preserving descriptive quality.
  • CHAIR Rohrbach et al. (2018) (Caption Hallucination Assessment with Image Relevance) quantifies hallucinations by detecting whether the model-generated captions mention objects that do not exist in the image. It provides two variants:
    $\text{CHAIR}_{I}=\frac{|\{\text{hallucinated objects}\}|}{|\{\text{all objects}\}|}$ (7)
    $\text{CHAIR}_{S}=\frac{|\{\text{hallucinated responses}\}|}{|\{\text{all responses}\}|}$ (8)
    where CHAIRI measures instance-level hallucination (object granularity) and CHAIRS measures sentence-level hallucination (response granularity).
  • POPE Li et al. (2023b) (Polling-based Object Probing Evaluation) automates hallucination detection via instance-level object probing. It:
    • Segments objects in the image;
    • Asks the model about object existence and introduces distractor queries;
    • Computes metrics such as F1 score to measure detection precision.
    POPE offers direct insights into a model’s visual grounding capability through objective visual questioning.
A.3.2 Metrics on Cross-modal Faithfulness
On the side of Cross-modal Faithfulness, we adopt Faith and FaithS (Jing et al., 2024), which evaluate how well the generated text is grounded in the visual input. Faith focuses on overall alignment, while FaithS specifically checks whether statements are supported by the visual evidence in a token-level or segment-wise manner. In addition, we report the $F1$ score, a standard metric that captures the harmonic mean of precision and recall between the predicted and reference entities. In our context, it reflects how well the model identifies relevant visual content without fabricating or omitting essential elements, thus serving as a practical indicator of the model’s grounding precision and completeness.
  • F1 Score reflects the harmonic mean of precision and recall in detecting whether queried objects exist. High F1 indicates accurate recognition and rejection of hallucinated entities:
    $\text{F1}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (9)
  • Faith (Jing et al., 2024) measures the overall semantic alignment between image and response. It uses automated matching or human verification to assess whether the content is factually grounded in the image:
    $\text{Faith}=\frac{|\text{Aligned Statements}|}{|\text{Total Statements}|}$ (10)
  • FaithS (Jing et al., 2024) extends Faith to a finer granularity by evaluating the support of specific sentence segments or tokens using cross-modal supervision or saliency alignment:
    $\text{Faith}_{S}=\frac{|\text{Grounded Segments or Tokens}|}{|\text{Total Segments or Tokens}|}$ (11)
A.4 Evaluation Benchmark
Besides, to provide a fine-grained and multi-perspective assessment of ReLoop’s effectiveness in suppressing hallucinations and enhancing cross-modal faithfulness, we adopt four complementary benchmarks. AMBER (Wang et al., 2023) targets object-level hallucinations, while MMHal-B (Sun et al., 2023) and HallusionBench (Guan et al., 2024) assess errors in attributes, spatial relations, and perceptual consistency. MME (Fu et al., 2023) covers general multimodal capabilities such as OCR and counting. These benchmarks collectively evaluate generative and discriminative capabilities, entity grounding, perceptual consistency, and multimodal reasoning:
  • AMBER (Wang et al., 2023): An LLM-free multi-dimensional benchmark that diagnoses hallucinations in both generative and discriminative tasks. It explicitly tests object existence, attributes, and relations, allowing us to assess ReLoop’s object-level grounding fidelity, attribute correctness, and relational accuracy. This supports the evaluation of semantic precision in visual grounding.
  • MMHal-B (Sun et al., 2023): A benchmark built upon fact-augmented reinforcement learning (RLHF) that penalizes hallucinated attributes and spatial configurations. MMHal-B offers targeted diagnostics for hallucination suppression in factual and compositional dimensions, particularly assessing whether ReLoop can resist overgeneralization and maintain factual grounding under complex prompts.
  • HallusionBench (Guan et al., 2024): A benchmark that probes visual-linguistic robustness under ambiguous image-text settings. It emphasizes contextual grounding, requiring models to handle subtle visual cues and nuanced linguistic traps. HallusionBench evaluates ReLoop’s ability to maintain perceptual consistency and reject misleading contextual cues that typically trigger hallucinations.
  • MME Fu et al. (2023): A broad-spectrum benchmark measuring multimodal perception and cognition across 14 sub-tasks, including OCR, object counting, spatial reasoning, and commonsense grounding. MME validates whether ReLoop’s structured supervision translates into generalized improvements in visual understanding and multimodal reasoning, beyond hallucination mitigation.
Together, these benchmarks offer layered supervision signals from fine-grained object hallucination detection to holistic multimodal cognition, providing strong empirical evidence of ReLoop’s reliability across diverse real-world tasks.
Figure 6: Case Study: Comparison between MiniGPT-4 and ReLoop across four types of hallucination in open-ended VQA: Object, Attribute, Relation, and Event. ReLoop produces more accurate and grounded responses by aligning its outputs with both the visual evidence and the question semantics.
A.5 Baseline Implementation
To evaluate ReLoop’s generalizability and additive benefit, we compare it with three representative alignment-based hallucination mitigation strategies: LLaVA-RLHF Sun et al. (2023), HA-DPO Zhao et al. (2023) , and POVID Zhou et al. (2024). These baselines span a diverse range of supervision paradigms, from reinforcement learning to contrastive grounding. Importantly, all methods are applied on top of the same backbone (LLaVA-1.5) with consistent training configurations, ensuring fair comparison.
  • LLaVA-RLHF Sun et al. (2023) aligns responses to human preferences through reinforcement learning from human feedback. While effective for improving general fluency and tone, it does not explicitly penalize visual or factual inconsistencies.
  • HA-DPO Zhao et al. (2023) adopts hallucination-aware preference optimization by contrasting faithful versus hallucinated generations. This method introduces targeted loss signals during fine-tuning, encouraging the model to avoid semantically spurious content.
  • POVID Zhou et al. (2024) enhances visual grounding via perturbed image inputs, injecting contrastive visual signals to reduce reliance on textual priors and promote visual fidelity.
  • VCD Leng et al. (2023) is a training-free, inference-time method that contrasts output distributions from the original image and a visually distorted counterpart, down-weighting tokens supported mainly by unimodal/language priors and thereby mitigating object hallucinations without further tuning.
Results from both fine-grained hallucination metrics (Table 4) and benchmark-level evaluations (Table 6) demonstrate that ReLoop consistently outperforms all competing methods. These results validate ReLoop as a robust and generalizable framework capable of enhancing multimodal model performance beyond what is achievable by current alignment-based techniques alone.
Method Trainable Params Feedback Modules GPU Hours GPU Type Peak VRAM Notes
ReLoop (MiniGPT-4) M+SM{+}S ($\sim$1.2B) Frozen 3 A100-40G $\sim$26GB Fastest; ablation
ReLoop (LLaVA-1.5) M+SM{+}S ($\sim$13B) Frozen 6 A100-80G $\sim$48GB Core experiment. (Sec. 5.2)
RLHF finetuning  Whole model ($\sim$13B) End-to-end 20 A100-80G $\sim$70GB
DPO Whole model End-to-end 8–16 A100-80G >>50GB
Contrastive alignment Whole model Encoder fusion $\sim$10 A100-80G $\sim$32GB
Table 7: Cross-method training cost. ReLoop updates only $M$ and $S$ with frozen CFPs, reducing optimizer/memory footprint relative to end-to-end baselines.
Variant CFP Modules GPU Hours Peak Mem Train Latency POPE $\uparrow$
[…]
A.6 Training Cost Breakdown
We quantify ReLoop’s computational footprint and efficiency trade-offs from four angles: (i) cross-method training cost, (ii) overhead attribution via CFP-component ablations, (iii) convergence versus training steps, and (iv) per-epoch cost under a controlled regime. Unless otherwise stated, CFP modules are frozen (no gradient), and only the main model $M$ and the Semantic Aggregator $S$ are updated.
A.6.1 Cross-Method Training Cost
We compare ReLoop against representative alignment methods. As summarized in Table 7, ReLoop updates only $M$ and $S$ with frozen CFPs, substantially reducing optimizer/memory footprint compared to end-to-end baselines. Despite extra forward passes for CFPs, ReLoop avoids optimizer states for large frozen modules, yielding substantially lower GPU hours and memory than end-to-end RLHF/DPO. This design preserves training affordability while enabling closed-loop supervision.
A.6.2 Overhead Attribution via CFP Components
We attribute runtime and memory overheads to individual CFP branches on LLaVA-1.5, keeping data/backbone constant. Table 8 reports GPU hours, memory, latency, and POPE. As seen in Table 8, both language and visual CFPs contribute positively to hallucination suppression, with additive gains. The per-branch overhead is bounded and predictable. Since CFPs are optional at inference, the training-time penalty does not translate to deployment latency. Both language and visual CFPs contribute positively to hallucination suppression, with additive gains. The overhead is bounded and predictable per component. Since CFPs are optional at inference, the training-time penalty does not translate to deployment latency.
A.6.3 Convergence and Epoch-Level Cost
[…]
Table 9: Per-epoch cost under the controlled regime. ReLoop increases epoch time by $\sim$77% yet converges within comparable steps. CFPs are disabled at inference, so deployment latency remains unchanged.
[…]
0 (no Training) 0.0 77.2
1k 3.6 80.1 +2.9 0.81
2k 7.2 81.6 +1.5 0.21
3k 10.9 82.2 +0.6 0.08
Table 10: Diminishing returns with steps (LLaVA-1.5). >>90% of gains appear within 2k steps ($\sim$7 GPU hours).
A.7 Contrastive Augmentation Ablation
To evaluate whether ReLoop depends on the manually perturbed semantic negatives (Appendix A.2), we ablate this contrastive augmentation and train on the standard LLaVA-Instruct data only, keeping all other settings identical to the main experiments. Even without contrastive augmentation, ReLoop significantly improves over the LLaVA-1.5 baseline on all metrics (Table 11), demonstrating that the core closed-loop alignment mechanism is effective when trained purely on standard supervision data. The contrastive examples provide further refinement, most notably on AMBER and MMHal-B, but they are not essential for ReLoop to outperform existing alignment strategies.
[…]
LLaVA-1.5
(baseline) 73.9 1513 65.4 48.6
ReLoop
(w/o augmentation) 78.5 1452 67.1 50.1
Full
ReLoop 80.3 1505 68.9 52.3
Table 11: Ablation on contrastive augmentation (LLaVA-1.5 backbone). Even without augmentation, ReLoop improves over the base model across all benchmarks; augmentation yields additional gains.
Appendix B Case Study
We present a qualitative case study to analyze how ReLoop mitigates hallucination across four representative types:
  • Object Hallucination: The baseline model incorrectly asserts the presence of a "referee stand" which is not in the image. ReLoop corrects this by recognizing the absence of such an entity.
  • Attribute Hallucination: An animal is mislabeled as "dog" instead of "chihuahua." ReLoop identifies the finer-grained attribute correctly.
  • Relation Hallucination: The spatial relationship "on a sofa" is incorrectly predicted; ReLoop grounds the child’s location more accurately.
  • Event Hallucination: The action "not playing" contradicts visual evidence; ReLoop revises the answer to match the depicted motion.
As shown in Figure 6, baseline models such as MiniGPT-4 frequently produce fluent yet inaccurate answers that are not grounded in the image. ReLoop corrects these errors by leveraging consistency feedback to align its answers with both the question intent and visual content. The examples highlight ReLoop’s capacity to suppress diverse hallucination patterns and improve factual reliability in open-ended VQA.
B.1 Handling Nonsensical or Unrelated Initial Answers
This section complements the above qualitative cases by focusing on an orthogonal failure mode raised by reviewers: how ReLoop handles instances where the initial answer $A$ is nonsensical or unrelated to the question. ReLoop employs a set of structured safeguards: early rejection, $\gamma$-based downweighting via ACW, and entropy-aware masking of attention. The set of structured safeguards can prevent misleading updates when the feedback is deemed unreliable. The representative examples in Table 12 illustrate how these filters operate at the case level before gradients are applied. In addition, the quantitative stress tests in the main text (see Table 5). demonstrate stable performance under injected noise, corroborating the robustness of these safeguards.
Failure Type Q A Action Taken
Empty Output “What is on the table?” (empty) Early rejection: sample skipped; no feedback loss computed.
Overly Generic “What sport is being played?” “I’m not sure.” $\gamma$-downweighting: low-confidence ACW weight; skipped if below length threshold.
Nonsensical Repetition “What is the man holding?” “Banana banana banana sky help!” Syntactic abnormality detection: reject via regex/#token heuristics; no update.
Unrelated Semantic “What color is the bus?” “Apples grow in the summer.” Semantic mismatch: $\gamma\!\approx\!0$ from BERTScore $\Rightarrow$ loss suppressed.
Hallucinated Words “What are the people doing?” “Grockling spinners do fleeb!” Entropy + token validation: flat attention masked; spurious tokens filtered before loss.
Table 12: Case-level handling of nonsensical or unrelated initial answers in ReLoop. Structured safeguards (early rejection, ACW $\gamma$-based downweighting, and entropy-aware masking) prevent misleading gradients from invalid feedback.
Appendix C Entropy-based Pseudo Ground-Truth Attention $\mathcal{H}_{\text{pseudo}}$
This section details the construction of the entropy-based pseudo ground-truth attention used in Sec. 4.2.3. We describe tensor shapes, multi-layer/head aggregation, token filtering, smoothing/normalization, and edge cases to facilitate reproduction.
Notation and Shapes.
Let the decoder cross-attention at decoding step $t$ be $\{A^{(\ell,h)}_{t}\in\mathbb{R}^{S}\}_{\ell=1..L,\,h=1..H}$ over $S$ visual patches (keys), for $L$ layers and $H$ heads. We first aggregate heads and layers to obtain a single distribution over patches for token $t$:
$\bar{a}_{t}\;=\;\sum_{\ell=1}^{L}\sum_{h=1}^{H}w^{(\ell)}\,u^{(h)}\,A^{(\ell,h)}_{t}$ (12)
$\sum_{\ell}w^{(\ell)}=\sum_{h}u^{(h)}=1$ (13)
where $w^{(\ell)}$ and $u^{(h)}$ are fixed convex weights. In a default setting, we set uniform across heads ($u^{(h)}{=}1/H$) and a back-loaded layer prior ($w^{(\ell)}\propto\exp(\kappa\,\ell/L)$ with $\kappa{=}1.5$) to emphasize later layers. We row-normalize to obtain a probability over spatial patches:
$p_{t,s}\;=\;\frac{\bar{a}_{t}[s]}{\sum_{s^{\prime}}\bar{a}_{t}[s^{\prime}]}\;\in\;[0,1],\qquad\sum_{s}p_{t,s}=1.$ (14)
Per-token Entropy and Confident Set.
For each generated token $t$, we excludes BOS/EOS/padding/special tokens and computes entropy:
$\mathcal{E}_{t}\;=\;-\sum_{s=1}^{S}p_{t,s}\log p_{t,s}\;\;\in\;[0,\log S]$ (15)
and form a confident token set:
$\mathcal{T}_{\text{conf}}\;=\;\{\,t\;\mid\;\mathcal{E}_{t}\leq\tau\,\}$ (16)
In a default setting, we set $\tau{=}2.0$ (nats). In low-entropy regimes ($\mathcal{E}_{t}\approx 0$) the attention is highly focused; high entropy indicates diffuse/unstable focus. In an optional variant (ablation only) setting, we reweight votes by $w_{t}{=}\max\!\big(0,1-\frac{\mathcal{E}_{t}}{\log S}\big)$; we do not apply this by default to keep the scheme simple and robust.
Voting and Temperature-normalized Map.
Confident tokens vote for patch importance by summation:
$\tilde{h}[s]\;=\;\sum_{t\in\mathcal{T}_{\text{conf}}}p_{t,s}$ (17)
We convert $\tilde{h}$ to a soft target by temperature softmax (Default: $T_{a}{=}0.7$):
$\mathcal{H}_{\text{pseudo}}[s]\;=\;\frac{\exp\!\big(\tilde{h}[s]/T_{a}\big)}{\sum_{s^{\prime}}\exp\!\big(\tilde{h}[s^{\prime}]/T_{a}\big)}$ (18)
Spatial Reshaping and Smoothing.
Let the $S$ patches correspond to a $(H_{p}\times W_{p})$ grid from the vision encoder (e.g., ViT patch tokens). We reshape $\mathcal{H}_{\text{pseudo}}\!\in\!\mathbb{R}^{S}$ to $\mathbb{R}^{H_{p}\times W_{p}}$, apply light Gaussian smoothing ($3{\times}3$ kernel, $\sigma{=}0.8$; reflect padding), then flatten back to length $S$. A final $\ell_{1}$ normalization ensures $\sum_{s}\mathcal{H}_{\text{pseudo}}[s]{=}1$.
Special Tokens, Padding, and Masking.
We exclude special tokens (BOS/EOS, padding) and punctuation-only tokens from $\mathcal{T}_{\text{conf}}$. For subword tokenization, all subpieces are treated uniformly; no external POS/saliency tools are used to preserve the unsupervised nature.
Empty-set and Degenerate Cases.
If $\mathcal{T}_{\text{conf}}=\varnothing$ (rare; e.g., extremely diffuse attention), we fall back to a min-entropy top-$k$ strategy: pick the $k{=}\max(1,\lceil 0.01T\rceil)$ lowest-entropy tokens to form $\mathcal{T}_{\text{conf}}$ and proceed with Eq. (17). If $\tilde{h}$ is flat (numerically), we use a near-uniform prior slightly peaked at the global min-entropy token’s argmax.
Algorithm 1 Constructing Entropy-based Pseudo Attention $\mathcal{H}_{\text{pseudo}}$
1:Cross-attn $\{A_{t}^{(\ell,h)}\!\in\!\mathbb{R}^{S}\}$, weights $\{w^{(\ell)}\}$, $\{u^{(h)}\}$, threshold $\tau$, temperature $T_{a}$
[…]
4:for each token $t$ in generated tokens (exclude specials) do
5:  $\bar{a}\!\leftarrow\!\sum_{\ell=1}^{L}\sum_{h=1}^{H}w^{(\ell)}u^{(h)}A_{t}^{(\ell,h)}$ $\triangleright$ Eq. 13
6:  $p\!\leftarrow\!\bar{a}/\sum_{s}\bar{a}[s]$ $\triangleright$ Eq. 14
7:  $\mathcal{E}_{t}\!\leftarrow\!-\sum_{s}p[s]\log p[s]$ $\triangleright$ Eq. 15
8:  if $\mathcal{E}_{t}\leq\tau$ then
9:   $\mathcal{T}_{\mathrm{conf}}\!\leftarrow\!\mathcal{T}_{\mathrm{conf}}\cup\{(t,p)\}$
10:  end if
11:end for
12:if $\mathcal{T}_{\mathrm{conf}}=\varnothing$ then
13:  $\mathcal{T}_{\mathrm{conf}}\!\leftarrow$ top-$k$ lowest-entropy tokens
14:end if
[…]
19:return $\mathrm{stopgrad}(\mathcal{H}_{\text{pseudo}})$
Objective and Gradient Flow.
The attention supervision minimizes a KL divergence between the model’s cross-attention map $\mathcal{H}$ (from the current forward pass) and $\mathcal{H}_{\text{pseudo}}$:
$\mathcal{L}_{\text{attn}}$ $=\mathrm{KL}\!\big(\mathcal{H}\,\|\,\mathcal{H}_{\text{pseudo}}\big)$ (19)
$=\sum_{s}\mathcal{H}[s]\Big(\log\mathcal{H}[s]-\log\mathcal{H}_{\text{pseudo}}[s]\Big)$
We stop gradients through $\mathcal{H}_{\text{pseudo}}$; only the main model’s attention is updated. The loss is weighted by $\gamma$ in the total objective (Sec. 4.2.1).
Pseudocode
The procedure (Algorithm  1) is unsupervised and self-adaptive, relying solely on model-internal attentional confidence and requiring no external saliency or human annotation. Its design (entropy gating, temperature control, and mild spatial smoothing) yields stable targets for the KL alignment used in Sec. 4.2.3.